OpenStudy (aryana_maria2323):

If f and g are differentiable functions for all real values of x such that f(1) = 4, g(1) = 3, f '(3) = −5, f '(1) = −4, g '(1) = −3, g '(3) = 2, then find h '(1) if h(x) = the quotient of f of x and g of x.. a.17 b. 11 c. 0 d. −11

OpenStudy (aryana_maria2323):

Can you please help? @imqwerty @tkhunny @wqfqlwekfwl @123AB456C

OpenStudy (jchick):

I just posted a link in the math chatroom so hopefully someone will help have a great day bye! :)

OpenStudy (aryana_maria2323):

Thank you and you have a wonderful day to! @jchick

OpenStudy (aryana_maria2323):

Can you please help me @IrishBoy123

OpenStudy (danjs):

hi

OpenStudy (aryana_maria2323):

can you please help me @DanJS

OpenStudy (danjs):

they define h(x) as \[\large h(x) = \frac{ f(x) }{ g(x) }\] take the derivative of both sides of that

OpenStudy (danjs):

they want h '(1)

OpenStudy (aryana_maria2323):

I don't know how to do that because the are two f'(x) and two g'(x)

OpenStudy (danjs):

have to use the quotient rule , take derivative of both sides of the h(x) function with respect to x

OpenStudy (aryana_maria2323):

How do I do that?

OpenStudy (aryana_maria2323):

Can you write it out so I can see the steps?

OpenStudy (danjs):

check this \[\large h '(x) = \frac{ g(x) * f '(x) - f(x)*g'(x)}{ [g(x)]^2 }\]

OpenStudy (aryana_maria2323):

Okay but there is two f'(x) and two g'(x). So which ones do I use?

OpenStudy (danjs):

they gave you both function values and both derivative values at x=1, just put those values in to get the h'(1) value

OpenStudy (danjs):

yeah they want you to find h'(1) , x=1

OpenStudy (aryana_maria2323):

so replace 1 for x in the entire problem?

OpenStudy (danjs):

use x=1 in h'(x) .... \[\large h '(1) = \frac{ g(1) * f '(1) - f(1)*g'(1)}{ [g(1)]^2 }\]

OpenStudy (danjs):

all those values are given , fill them in

OpenStudy (aryana_maria2323):

okay that is exactly what I got

OpenStudy (aryana_maria2323):

so then replace g and f with the numbers given in the problem?

OpenStudy (danjs):

yes, all those are given

OpenStudy (danjs):

g(1) f(1) g ' (1) f ' (1)

OpenStudy (danjs):

you understand it all now or no

OpenStudy (aryana_maria2323):

yes I am trying to work it out. one second.

OpenStudy (aryana_maria2323):

 is this correct?